A body is projected with velocity of 20mps in a direction making angle of 60 degree with horizontal . Calculate (1) position after 0.5 sec and  (2) velocity after 0.5 sec .

Grade:11

1 Answers

Anitha Rajkumar
6 years ago
Let the angle be cX=(ucosc)t =20×0.5×0.5 =5mY=(usinc)t-1/2gt^2 =20×root3/2 ×0.5 - 1/2 ×10×0.25 =7.4mResultant of x and y will be root of x^2+y^2= Root of 70.76 =8.94m